OpenStudy (anonymous):

Add the following base 2 numbers 1000101 and 10011

OpenStudy (campbell_st):

remember with base 2 numbers there are only 2 digits 0 and 1 so 1 + 1 = 10 in base 2 1000101 (69) 10011 (19) --------- 1011000 (88) and to check \[88 = 1 \times 2^6 + 0 \times 2^5 + 1 \times 2^4 + 1 \times 2^3 + 0 \times 2^2 + 0 \times 2^1 + 0 \times 2^0\] hope it helps